# Magellan: Guided MCTS for Latent Space Exploration and Novelty Generation

### v1

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable proficiency across a wide array of complex generation tasks, establishing new frontiers in fields ranging from creative writing to scientific discovery Brown et al. (2020); Zhao et al. (2023). However, the very autoregressive nature that underpins their fluency also renders them susceptible to settling into the “gravity wells” of their own knowledge. This phenomenon stems directly from their training objective of maximizing local conditional probabilities, which inherently biases generation towards high-density regions within the training data’s distribution. Consequently, these models excel at producing ‘safe’, high-frequency sequences that mimic familiar patterns but often fail to explore the long tail of the distribution where novel ideas reside. This myopic, greedy strategy becomes a critical bottleneck for tasks requiring long-range planning or the creative synthesis of disparate concepts. When faced with such challenges, LLMs tend to converge prematurely into sub-optimal modes, struggling to discover globally optimal or truly innovative solutions.
In an effort to steer generation away from these probabilistic gravity wells, a variety of inference-time “guidance tools” have been developed, which can be broadly categorized by their level of intervention. The most direct approaches operate at the decoding level. Stochastic sampling methods, such as Top-p Holtzman et al. (2019), inject randomness to diversify outputs, but this often results in an unguided random walk through the semantic space that sacrifices coherence. Search-based decoding algorithms like Beam Search Graves (2012), while less random, attempt to approximate a global search but are constrained by a fixed beam width, frequently pruning high-potential, non-obvious pathways too early. A more structured class of methods operates at the cognitive level via prompt engineering. Chain-of-Thought (CoT) prompting Wei et al. (2022), for instance, guides the model through a pre-defined, linear reasoning process. This represents a form of static planning; the reasoning path is fixed and cannot adapt to challenges encountered during generation. The most advanced baselines embrace a fully dynamic search paradigm. Frameworks like Tree of Thoughts (ToT) Yao et al. (2023a) explicitly model generation as exploring a tree of possibilities. This is a powerful conceptual leap, yet these frameworks critically falter on their evaluation mechanism. They typically rely on the LLM to self-evaluate the quality of different branches, a process of self-reflection that is often unprincipled, inconsistent, and lacks a clear objective function. This exposes a crucial research gap: the need for a framework that not only searches dynamically but is also guided by a principled, explicit, and multi-objective evaluation strategy, steering the exploration towards a globally desirable outcome rather than merely a locally plausible one.
@@ 2.1 LLMs for Scientific Discovery and Structured Reasoning
The automation of scientific discovery, particularly for novel idea generation, has become a key application for LLMs. One major research thrust synthesizes new concepts by retrieving and recombining facets of existing literature Wang et al. (2024); Radensky et al. (2024); Li et al. (2024). However, these approaches often rely on the LLM’s textual understanding and splicing, lacking the underlying control needed for precise conceptual navigation. Another direction mimics scientific collaboration using multi-agent systems to brainstorm and refine ideas Su et al. (2024); Ghafarollahi and Buehler (2025). The viability of LLMs as innovators is underscored by large-scale human studies Si et al. (2024) and the creation of dedicated benchmarks Guo et al. (2025). To tackle such complex, multi-step tasks, the community has developed structured reasoning frameworks. These have evolved from linear Chain-of-Thought (CoT) prompting Wei et al. (2022) and tool-augmented reasoning Yao et al. (2023b) to more dynamic search over a Tree or Graph of Thoughts (ToT/GoT)  Yao et al. (2023a); Besta et al. (2024). However, the efficacy of these search-based methods is often constrained by simple heuristics or a reliance on the LLM’s own inconsistent self-evaluation to guide exploration  Feng et al. (2025), a limitation our work directly addresses.
@@ 2.2 Advanced Search for Novelty and Diversity
A central challenge in generative AI is ensuring the novelty and diversity of outputs, a problem quantified by benchmarks like NoveltyBench Zhang et al. (2025). Various techniques aim to enhance novelty. Some diversify prompts via multi-view embeddings Lagzian et al. , which, while providing multiple entry points, remain constrained by the LLM’s existing knowledge manifold and struggle to escape high-probability conceptual regions. Others explore the model’s latent space through simple interpolation Bystroński et al. (2025), but this primarily navigates between known points rather than forging truly new conceptual paths. The assessment of novelty itself also remains an active research area Lin et al. (2024); Liu et al. (2025). Our work proposes a more foundational solution by incorporating Monte Carlo Tree Search (MCTS), a powerful algorithm from reinforcement learning renowned for its ability to balance the exploration-exploitation dilemma. MCTS has a strong track record in complex generative domains where strategic planning is paramount, including strategic synthesis Silver et al. (2016) and agentic reasoning Luo et al. (2025).By integrating MCTS with a principled, multi-objective value function, Magellan moves beyond simple diversification to perform a strategic, goal-directed search for globally innovative solutions, addressing the evaluation gap in prior search-based frameworks.
@@ 4.1 Experimental Setup
To ground our model, we constructed a knowledge corpus of 16,582 paper abstracts from top-tier venues (e.g., CVPR, ICML, Nature Medicine). We used Qwen3-1.7B Yang et al. (2025) to encode them into a FAISS-indexed Douze et al. (2024) vector database for efficient similarity search. To ensure a fair and challenging evaluation, we generated a set of 200 diverse, cross-disciplinary research themes as a theme pool, and randomly sampled 50 themes to form the core test set . Further details on corpus construction and theme generation are provided in the appendix.
@@
We compare Magellan against a suite of strong reasoning and generation baselines: Zero-shot Brown et al. (2020), Chain-of-Thought (CoT) Wei et al. (2022), ReAct Yao et al. (2023b), and Tree of Thoughts (ToT) Yao et al. (2023a).
@@
For a fair comparison, all methods used Qwen3-1.7B as the backbone model. Magellan’s MCTS was configured with a maximum of 30 iterations, but our efficiency mechanisms (see Sec. 3) led to an average of only $\sim$3 iterations before convergence. We employed an LLM-as-a-Judge protocol using DeepSeek-V3.1-Think model Liu et al. (2024) as the impartial evaluator. The judge scored each output on a 1-10 scale for Plausibility, Clarity, and Innovation. It also made a forced-choice selection for the Overall Best proposal, from which we calculate the Win Rate. This rate reflects the outcome of a direct, forced-choice comparison within each specific experiment, and is therefore relative to the opponents in that comparison. All experiments were conducted on a four-node Tesla V100 GPU server, with a total experiment time of approximately 36 hours.
@@ 4.3 Comparison with AI-for-Science Frameworks
We benchmarked Magellan against two SOTA AI-for-Science frameworks, AI Scientist Lu et al. (2024) and SciPip Wang et al. (2024), on the core task of idea expansion. For a fair comparison, all methods used the same initial ideas and Qwen3-1.7B backbone. We adapted the baselines to focus only on idea expansion: for AI Scientist, we used its generation module prompt; for SciPip, we used its expansion component directly. To align the output from AI Scientist with the depth expected in our task, we augmented its prompt to require a detailed elaboration paragraph. The final evaluated text combined this elaboration with the generated experimental steps.
@@ A.3 Implementation and Hyperparameter Details
Across all experiments, for both our method and the baselines, the core Large Language Model was Qwen3-1.7B Yang et al. [2025]. For text generation, we consistently used a sampling temperature of 0.7 and a top-p value of 0.9 to encourage creative yet coherent outputs.
@@ A.4 LLM-as-a-Judge Protocol
To ensure a transparent and reproducible evaluation process, this section details the protocol used for our LLM-as-a-Judge methodology. We utilized the DeepSeek-V3.1-Think model Liu et al. [2024] as the core evaluator. To mitigate potential bias, the proposals from all methods were presented to the judge in a randomized order and without any identifying labels. The evaluation was guided by a prompt shown in A.4.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable proficiency across a wide array of complex generation tasks, establishing new frontiers in fields ranging from creative writing to scientific discovery brown2020language; zhao2023survey. However, the very autoregressive nature that underpins their fluency also renders them susceptible to settling into the “gravity wells” of their own knowledge. This phenomenon stems directly from their training objective of maximizing local conditional probabilities, which inherently biases generation towards high-density regions within the training data’s distribution. Consequently, these models excel at producing ‘safe’, high-frequency sequences that mimic familiar patterns but often fail to explore the long tail of the distribution where novel ideas reside. This myopic, greedy strategy becomes a critical bottleneck for tasks requiring long-range planning or the creative synthesis of disparate concepts. When faced with such challenges, LLMs tend to converge prematurely into sub-optimal modes, struggling to discover globally optimal or truly innovative solutions.
In an effort to steer generation away from these probabilistic gravity wells, a variety of inference-time “guidance tools” have been developed, which can be broadly categorized by their level of intervention. The most direct approaches operate at the decoding level. Stochastic sampling methods, such as Top-p holtzman2020curious, inject randomness to diversify outputs, but this often results in an unguided random walk through the semantic space that sacrifices coherence. Search-based decoding algorithms like Beam Search graves2012sequence, while less random, attempt to approximate a global search but are constrained by a fixed beam width, frequently pruning high-potential, non-obvious pathways too early. A more structured class of methods operates at the cognitive level via prompt engineering. Chain-of-Thought (CoT) prompting wei2022chain, for instance, guides the model through a pre-defined, linear reasoning process. This represents a form of static planning; the reasoning path is fixed and cannot adapt to challenges encountered during generation. The most advanced baselines embrace a fully dynamic search paradigm. Frameworks like Tree of Thoughts (ToT) yao2023tree explicitly model generation as exploring a tree of possibilities. This is a powerful conceptual leap, yet these frameworks critically falter on their evaluation mechanism. They typically rely on the LLM to self-evaluate the quality of different branches, a process of self-reflection that is often unprincipled, inconsistent, and lacks a clear objective function. This exposes a crucial research gap: the need for a framework that not only searches dynamically but is also guided by a principled, explicit, and multi-objective evaluation strategy, steering the exploration towards a globally desirable outcome rather than merely a locally plausible one.
@@ 2.1 LLMs for Scientific Discovery and Structured Reasoning
The automation of scientific discovery, particularly for novel idea generation, has become a key application for LLMs. One major research thrust synthesizes new concepts by retrieving and recombining facets of existing literature wang2024scipip; chan2024scideator; li2024chain. However, these approaches often rely on the LLM’s textual understanding and splicing, lacking the underlying control needed for precise conceptual navigation. Another direction mimics scientific collaboration using multi-agent systems to brainstorm and refine ideas su2024many; ai2024sciagents. The viability of LLMs as innovators is underscored by large-scale human studies si2024can and the creation of dedicated benchmarks guo2024ideabench. To tackle such complex, multi-step tasks, the community has developed structured reasoning frameworks. These have evolved from linear Chain-of-Thought (CoT) prompting wei2022chain and tool-augmented reasoning yao2023react to more dynamic search over a Tree or Graph of Thoughts (ToT/GoT)  yao2023tree; besta2024graph. However, the efficacy of these search-based methods is often constrained by simple heuristics or a reliance on the LLM’s own inconsistent self-evaluation to guide exploration  zhang2024grapheval, a limitation our work directly addresses.
@@ 2.2 Advanced Search for Novelty and Diversity
A central challenge in generative AI is ensuring the novelty and diversity of outputs, a problem quantified by benchmarks like NoveltyBench zhang2024noveltybench. Various techniques aim to enhance novelty. Some diversify prompts via multi-view embeddings holm2024enhancing, which, while providing multiple entry points, remain constrained by the LLM’s existing knowledge manifold and struggle to escape high-probability conceptual regions. Others explore the model’s latent space through simple interpolation gao2024large, but this primarily navigates between known points rather than forging truly new conceptual paths. The assessment of novelty itself also remains an active research area lin2024evaluating; liu2024harnessing. Our work proposes a more foundational solution by incorporating Monte Carlo Tree Search (MCTS), a powerful algorithm from reinforcement learning renowned for its ability to balance the exploration-exploitation dilemma. MCTS has a strong track record in complex generative domains where strategic planning is paramount, including strategic synthesis silver2016mastering and agentic reasoning luo2025kbqa.By integrating MCTS with a principled, multi-objective value function, Magellan moves beyond simple diversification to perform a strategic, goal-directed search for globally innovative solutions, addressing the evaluation gap in prior search-based frameworks.
@@ 4.1 Experimental Setup
To ground our model, we constructed a knowledge corpus of 16,582 paper abstracts from top-tier venues (e.g., CVPR, ICML, Nature Medicine). We used Qwen3-1.7B qwen3 to encode them into a FAISS-indexed douze2024faiss vector database for efficient similarity search. To ensure a fair and challenging evaluation, we generated a set of 200 diverse, cross-disciplinary research themes as a theme pool, and randomly sampled 50 themes to form the core test set . Further details on corpus construction and theme generation are provided in the appendix.
@@
We compare Magellan against a suite of strong reasoning and generation baselines: Zero-shot brown2020language, Chain-of-Thought (CoT) wei2022chain, ReAct yao2023react, and Tree of Thoughts (ToT) yao2023tree.
@@
For a fair comparison, all methods used Qwen3-1.7B as the backbone model. Magellan’s MCTS was configured with a maximum of 30 iterations, but our efficiency mechanisms (see Sec. 3) led to an average of only $\sim$3 iterations before convergence. We employed an LLM-as-a-Judge protocol using DeepSeek-V3.1-Think model deepseek_placeholder as the impartial evaluator. The judge scored each output on a 1-10 scale for Plausibility, Clarity, and Innovation. It also made a forced-choice selection for the Overall Best proposal, from which we calculate the Win Rate. This rate reflects the outcome of a direct, forced-choice comparison within each specific experiment, and is therefore relative to the opponents in that comparison. All experiments were conducted on a four-node Tesla V100 GPU server, with a total experiment time of approximately 36 hours.
@@ 4.3 Comparison with AI-for-Science Frameworks
We benchmarked Magellan against two SOTA AI-for-Science frameworks, AI Scientist lu2024ai and SciPip wang2024scipip, on the core task of idea expansion. For a fair comparison, all methods used the same initial ideas and Qwen3-1.7B backbone. We adapted the baselines to focus only on idea expansion: for AI Scientist, we used its generation module prompt; for SciPip, we used its expansion component directly. To align the output from AI Scientist with the depth expected in our task, we augmented its prompt to require a detailed elaboration paragraph. The final evaluated text combined this elaboration with the generated experimental steps.
@@ 5 Discussion and Conclusion
Reproducibility Statement
Reproducing the Final Experimental Results
To ensure full reproducibility of our findings, our supplementary package contains all necessary components. Specifically, it includes: (1) Source Code: The complete source code for the Magellan framework and all evaluation scripts. (2) Data: The data required to replicate our experiments, including the identifiers for the knowledge corpus and the test set of themes. (3) Configuration: Configuration files detailing all hyperparameters and a requirements.txt file for the software environment.
Reproducing the AI-driven Discovery Process
The primary goal of our supplementary package is to make the agent-driven research process transparent, understandable, and conceptually reproducible. Transparency: The AgentChatLog/ directory provides a complete, step-by-step record of the prompts, agent outputs, and intermediate artifacts. This serves as a "lab notebook" for the AI’s work. Conceptual Reproducibility: While the stochastic nature of large language models means that identical outputs cannot be guaranteed, another researcher can follow the dialogues to understand the methodology. By using similar prompts and guidance, one could reasonably expect to guide a comparable AI agent to a similar set of ideas, code, and scientific conclusions. The logs document the critical path of inquiry, including dead ends, refinements, and breakthroughs. We believe this detailed record is essential for evaluating the methodology of using agents in science and provides a strong foundation for other researchers to build upon our approach.
Responsible AI Statement
Magellan is designed to augment human creativity and accelerate scientific discovery. We recognize that any powerful generative tool carries risks, such as generating plausible-sounding misinformation or amplifying biases from its training data. Therefore, we strongly assert that Magellan is a tool for assisting experts, not a replacement for them. Human oversight and expert validation of all outputs are the most critical safety measures for its responsible deployment. Our framework’s design incorporates further precautions by grounding the search in a peer-reviewed knowledge corpus and by providing full transparency through our open-source code. We believe that when used under expert supervision, Magellan can serve as a safe and powerful partner for innovation.
AI Agent Setup
The AI agent used in this research was configured within the gemini-cli interactive environment, which utilized the Gemini 2.5 Pro large language model as its core reasoning engine. The orchestration was managed by the gemini-cli framework, a stateful system that enables the LLM to decompose high-level objectives into sequential, executable tasks. This setup provided the agent with direct access to a suite of tools for interacting with the local development environment. Key tool integrations included: a comprehensive set of functions for file system manipulation (read_file, write_file, glob, search_file_content); and the ability to execute shell commands (run_shell_command) for compiling code and running experiments.
@@ A.3 Implementation and Hyperparameter Details
Across all experiments, for both our method and the baselines, the core Large Language Model was Qwen3-1.7B qwen3. For text generation, we consistently used a sampling temperature of 0.7 and a top-p value of 0.9 to encourage creative yet coherent outputs.
@@ A.4 LLM-as-a-Judge Protocol
To ensure a transparent and reproducible evaluation process, this section details the protocol used for our LLM-as-a-Judge methodology. We utilized the DeepSeek-V3.1-Think model deepseek_placeholder as the core evaluator. To mitigate potential bias, the proposals from all methods were presented to the judge in a randomized order and without any identifying labels. The evaluation was guided by a prompt shown in A.4.